Given that a positive charge Q1=+3c and negative charge Q2=2c, are separated by a distance of 8cm. Take permittivity constant k=9.0×10^9NM^2/c^2.

(a). find The electric field at point p 1cm from Q2

(b) what is the electric potential at point P ,5 cm from Q1.



1
Expert's answer
2021-06-01T10:50:36-0400

(a) The electric field at point P 1cm from Q2 can be found as follows:


"E=E_1-E_2=k(\dfrac{Q_1}{(x+d)^2}-\dfrac{Q_2}{x^2}),""E=9\cdot10^9\ \dfrac{N\cdot m^2}{C^2}\cdot(\dfrac{3\cdot10^{-6}\ C}{(0.01\ m+0.08\ m)^2}-\dfrac{2\cdot10^{-6}\ C}{(0.01\ m)^2}),""E=-1.76\cdot10^{8}\ \dfrac{N}{C}."

The sign minus means that the net electric field at point P directed leftward (towards the charge Q1).

(b) The electric potential at point P 5cm from Q1 can be found as follows:


"V=V_1+V_2,""V=\dfrac{kQ_1}{x}+\dfrac{kQ_2}{d+x}=k(\dfrac{Q_1}{x}+\dfrac{Q_2}{d+x}),""V=9\cdot10^9\ \dfrac{N\cdot m^2}{C^2}\cdot(\dfrac{3\cdot10^{-6}\ C}{0.05\ m}+\dfrac{-2\cdot10^{-6}\ C}{0.05\ m+0.08\ m}),""V=4\cdot10^5\ V."